OpenStudy (anonymous):

factor completely z^3-16

hartnn (hartnn):

are u sure its z^3-16 ? isn't it x^4-16 or z^2-16 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh no it says z^3-16z

hartnn (hartnn):

this makes it very easy! z^3-16z factor out the 'z' , what do u get ?

OpenStudy (anonymous):

(z2-4) (z+4z) ?

hartnn (hartnn):

no...how u got that ? and take one step at a time, factoring out z from z^3-16z will give z (z^2-16) got this? now can u factor z^2-16 ?

OpenStudy (anonymous):

no but i need to factor it out so i get a 2 binomials that when you multiply them equals z^3-16z

hartnn (hartnn):

i got the question, i asked u whether u got how, z^3-16z=z (z^2-16)

hartnn (hartnn):

\(\large z^2-16x = z^2\times z-16 \times z= z \times (z^2-16)\) yes?

OpenStudy (anonymous):

like for example i also had w^4-81 and the correct answer for that was (w^2-9) (w^2+9)

hartnn (hartnn):

yes, thats correct, but you first need to get the form \(\Large a^2-b^2\) to write it as \(\huge (a+b)(a-b)\)

hartnn (hartnn):

and z^2-16 is in that form, right ? (16=4^2)